### v1

# An Ising Similarity Regression Model

Abstract: Understanding the dependence structure between response variables is an important component in the analysis of correlated multivariate data. This article focuses on modeling dependence structures in multivariate binary data, motivated by a study aiming to understand how patterns in different U.S. senators’ votes are determined by similarities (or lack thereof) in their attributes, e.g., political parties and social network profiles. To address such a research question, we propose a new Ising similarity regression model which regresses pairwise interaction coefficients in the Ising model against a set of similarity measures available/constructed from covariates. Model selection approaches are further developed through regularizing the pseudo-likelihood function with an adaptive lasso penalty to enable the selection of relevant similarity measures. We establish estimation and selection consistency of the proposed estimator under a general setting where the number of similarity measures and responses tend to infinity. Simulation study demonstrates the strong finite sample performance of the proposed estimator in terms of parameter estimation and similarity selection. Applying the Ising similarity regression model to a dataset of roll call voting records of 100 U.S. senators, we are able to quantify how similarities in senators’ parties, businessman occupations and social network profiles drive their voting associations. Key words and phrases: Conditional dependence, Ising model, Lasso, Model selection, Multivariate data, Pseudo-likelihood
1 Introduction
The study of the dependence structure in correlated multivariate data has drawn much attention in recent years, due to the increasing accessibility of multi-response data across many disciplines such as finance (Tsay, 2013), economics (McElroy and Trimbur, 2023) and ecology (Hui, Müller, and Welsh, 2023). For a multivariate response vector $\bm{y}=(y_{1},\cdots,y_{p})^{\top}$, the dependencies across any set of two components $y_{j}$ and $y_{j^{\prime}}$ for $j,j^{\prime}=1,\cdots,p$ can be modeled by specifying a joint probabilistic distribution for $\bm{y}$. Two prominent examples are Gaussian graphical models (Whittaker, 1990) and Ising models (Ising, 1925), which specify a probability density function (pdf) and a probability mass function (pmf) for a continuous and binary response vector $\bm{y}$, respectively, where the pdf and pmf of these two models both contain a component of the form $\sum_{1\leq j<j^{\prime}\leq p}\exp(\theta_{jj^{\prime}}y_{j}y_{j^{\prime}})$ with $\theta_{jj^{\prime}}$ denoting a coefficient for the interaction term between $y_{j}$ and $y_{j^{\prime}}$. Accordingly, the set of $\theta_{jj^{\prime}};j,j^{\prime}=1,\cdots,p$} are referred to as pairwise interaction coefficients in these joint models. Moreover, in both models each $\theta_{jj^{\prime}}$ is directly tied to the conditional dependence relationship between $y_{j}$ and $y_{j^{\prime}}$ given the remaining elements in the response vector $\bm{y}$ (Hammersley-Clifford equivalence, Hammersley and Clifford, 1971): if $\theta_{jj^{\prime}}$ is close to/far away from zero, the conditional dependence between $y_{j}$ and $y_{j^{\prime}}$ is weak/strong, with the limiting case of $\theta_{jj^{\prime}}=0$ implying conditional independence. This interpretation of the pairwise interaction coefficient $\theta_{jj^{\prime}}$ has contributed to the popularity of Gaussian graphical models and Ising models for correlated multivariate data analysis; see for instance Meinshausen and Bühlmann (2006), Yuan and Lin (2007) and Friedman, Hastie, and Tibshirani (2007) for the Gaussian graphical models, and Majewski, Li, and Ott (2001), Guo et al. (2015) and Bhattacharya and Mukherjee (2018) for the Ising models, among many others. In this article, we focus on modeling the dependence relationships for a multivariate binary response vector $\bm{y}$ through, or equivalently, modeling the pairwise interaction coefficients $\theta_{jj^{\prime}}$ in, the Ising model. As demonstrated above, the conditional dependence of multivariate binary responses in the Ising model is captured by the pairwise interaction coefficients $\theta_{jj^{\prime}}$, which can be collected into a symmetric interaction matrix $\bm{\Theta}=(\theta_{jj^{\prime}})_{p\times p}$ where $\theta_{j^{\prime}j}=\theta_{jj^{\prime}}$. A variety of approaches have been proposed to estimate and regularize $\bm{\Theta}$. For example, Banerjee, El Ghaoui, and d’ Aspremont (2008) proposed a block-coordinate descent algorithm to solve an approximate sparse maximum likelihood problem for the estimation of $\bm{\Theta}$. Ravikumar, Wainwright, and Lafferty (2010) considered a neighborhood estimation method based on fitting a separate regularized logistic regression with the lasso penalty to each binary response, while Höfling and Tibshirani (2009) and Guo et al. (2010) used a pseudo-likelihood function coupled with a lasso-type penalty to simultaneously estimate and regularize all pairwise interaction coefficients in $\bm{\Theta}$. Further penalized likelihood approaches for sparse $\bm{\Theta}$ estimation have been developed by Lee, Ganapathi, and Koller (2006) and Xue, Zou, and Cai (2012), among others. In contrast to estimating and regularizing the elements of the interaction matrix $\bm{\Theta}$, this article proposes a novel Ising similarity regression model which regresses the pairwise interaction coefficients $\theta_{jj^{\prime}}$ against a set of pairwise similarity measures $w_{jj^{\prime}}^{(k)}$; that is,
$\theta_{jj^{\prime}}=\sum_{k=1}^{K}\alpha_{k}w_{jj^{\prime}}^{(k)},\textrm{ for }j\neq j^{\prime},$ (1.1)
where $\alpha_{k}$ denotes the regression coefficient associated with the $k$-th similarity measure $w_{jj^{\prime}}^{(k)}$ for $k=1,\cdots,K$. It is worth noting $w_{jj^{\prime}}^{(k)}$ measures the similarity between $j$ and $j^{\prime}$, which can be either observed directly as part of the data collection process, or induced from available auxiliary information. As a motivating example, we consider U.S. Senate roll call voting data where the binary response $y_{j}$ represents the $j$-th senator’s voting record (Yea or Nay) to a particular piece of legislation or bill. The pairwise interaction coefficient $\theta_{jj^{\prime}}$ thus captures the conditional dependence relationship between the $j$-th and $j^{\prime}$-th senators’ voting patterns, while $w_{jj^{\prime}}^{(k)}$ are constructed based on additional attributes of the $j$-th and $j^{\prime}$-th senators, such as political parties, ages, occupations and social network profiles. In such an example, the regression coefficients $\alpha_{k}$ in model (1.1) offer a clear, explicit quantification of how the $k$-th similarity measure between the $j$-th and $j^{\prime}$-th senators affects their conditional dependence when it comes to voting patterns.
At this point, it is important to acknowledge the work of Cheng et al. (2014), who proposed an Ising regression model to regress the pairwise interaction coefficients $\theta_{jj^{\prime}}$ on a covariate vector $\bm{x}$. Such a model with $\theta_{jj^{\prime}}=\theta_{jj^{\prime}}(\bm{x})$ does not provide the same interpretation for our motivating example however, as their interaction coefficients $\theta_{jj^{\prime}}$ for different pairs $(j,j^{\prime})$ only depend on the same covariate vector $\bm{x}$ instead of the pairwise similarity measures $w_{jj^{\prime}}^{(k)}$ considered in model (1.1).
Apart from the aforementioned example in political science, the proposed Ising similarity regression model has a multitude of applications across finance and ecology. For example, in the study of capital markets, $y_{j}$ can be an indicator denoting whether the $j$-th firm has distributed dividends to its shareholders, and $w_{jj^{\prime}}^{(k)}$ can be the similarity between the $j$-th and $j^{\prime}$-th firms’ financial fundamentals such as market value, cash flow, and leverage. Also, in ecology, $y_{j}$ can be a binary variable indicating the presence or absence of the $j$-th species, and $w_{jj^{\prime}}^{(k)}$ can be the similarity between the $j$-th and $j^{\prime}$-th species’ traits including but not limited to breeding season and overwintering behavior. For both of these examples, $\alpha_{k}$ in model (1.1) quantifies the effect of the $k$-th similarity measure on the conditional dependence relationships between the responses of the $j$-th and ${j^{\prime}}$-th companies or species. More broadly, the use of similarity measures in regression is also motivated by recent developments in covariance regression modeling (Zou et al., 2017, 2020, 2022), where the covariance between continuous responses is modeled as a linear combination of similarity measures. To the best of our knowledge though, the proposed Ising similarity regression model is the first to explicitly link similarity measures to the pairwise interaction coefficients (conditional dependence) of binary responses.
To estimate the Ising similarity regression model, we propose a regularized pseudo-likelihood approach which augments the pseudo-likelihood function of model (1.1) with an adaptive lasso penalty (Zou, 2006). Doing so induces sparse estimation of the regression coefficients $\{\alpha_{k};k=1,\cdots,K\}$, which allows us to recover similarity measures that are truly relevant in explaining the conditional dependence relationships between the binary responses. Under a setting where the number of regression coefficients $K$ and responses $p$ tend to infinity, we establish estimation and selection consistency for the regularized pseudo-likelihood estimator. To select the tuning parameter in the adaptive lasso penalty, we employ a cross-validation approach which preserves the dependencies between the elements of $\bm{y}$. Simulation results support the theoretical findings of the proposed estimator, demonstrating its strong finite sample estimation and model selection performance. We apply the Ising similarity regression model to roll call voting records of 100 U.S. senators from the 117-th Congress, with results demonstrating how similarities of senators’ attributes and social network activities drive the association between their voting patterns. In particular, aside from the expected findings such as senators from the same state or party being more likely to vote similarly, we find that senators who are businessmen or share certain follower-followee relationships on Twitter tend to exhibit more similar voting patterns.
The rest of this article is organized as follows. Section 2 introduces the Ising similarity regression model along with the proposed regularized pseudo-likelihood estimator. Section 3 discusses the theoretical properties of the regularized estimator. Section 4 presents simulation studies, while an application to U.S. roll call voting dataset is provided in Section 5. Section 6 offers some concluding remarks. All theoretical proofs of the theorems developed in this article are presented in the supplementary material.
2 An Ising Similarity Regression Model
2.1 Model Set-Up
Let $\bm{u}=(u_{1},\cdots,u_{p})^{\top}$ be any vector in the space $\{0,1\}^{p}$. The Ising model (Ising, 1925) specifies the following pmf for the $p$-dimensional binary response vector $\bm{y}$,
$f(\bm{u};\bm{\theta})=P(\bm{y}=\bm{u};\bm{\theta})=\frac{1}{Z(\bm{\theta})}\exp\left(\sum_{j=1}^{p}\theta_{jj}u_{j}+\sum_{1\leq j<j^{\prime}\leq p}\theta_{jj^{\prime}}u_{j}u_{j^{\prime}}\right),$ (2.2)
where $\bm{\theta}=(\theta_{11},\cdots,\theta_{1p},\cdots,\theta_{p-1,p-1},\theta_{p-1,p},\theta_{pp})^{\top}$ is a $p(p+1)/2$-dimensional parameter vector, and the partition function $Z(\bm{\theta})=\sum_{\bm{u}\in\{0,1\}^{p}}\exp(\sum_{j=1}^{p}\theta_{jj}u_{j}+\sum_{1\leq j<j^{\prime}\leq p}\theta_{jj^{\prime}}u_{j}u_{j^{\prime}})$ is an intractable normalization constant in the pmf. The main effect parameters in (2.2) are given by $\theta_{jj}$ for $j=1,\cdots,p$, while the pairwise interaction coefficients in the pmf are given by $\theta_{jj^{\prime}}$ for $j,j^{\prime}=1,\cdots,p$ and $\theta_{j^{\prime}j}=\theta_{jj^{\prime}}$. As discussed in Section 1, each of the pairwise interaction coefficients $\theta_{jj^{\prime}}$ is related to the conditional dependence between $y_{j}$ and $y_{j^{\prime}}$, conditional on the other elements of the response vector $\bm{y}$ (Hammersley and Clifford, 1971). This makes the Ising model useful in studying the conditional dependence structure among binary responses, through analyzing the symmetric interaction matrix $\bm{\Theta}=(\theta_{jj^{\prime}})_{p\times p}$.
Suppose now that, in addition to observing the binary response vector $\bm{y}$, we also record a set of $p\times p$ symmetric similarity matrices $\bm{W}_{k}=(w_{jj^{\prime}}^{(k)})_{p\times p}$ for $k=1,\cdots,K$, which may be available directly as part of the data collection process or constructed from auxiliary information variables $z_{j1},\cdots,z_{jK}$ associated to the $j$-th response for $j=1,\cdots,p$. For instance, in the motivating U.S. senators’ voting records data, one of the similarity matrices $\bm{W}_{k}$ is set as the $p\times p$ network adjacency matrix based on the follower-followee relationships on Twitter, while other similarity matrices are constructed based on auxiliary information variables $z_{jk}$ consisting of the $j$-th senator’s demographic attributes, such as political party, age, and occupation (see Section 5 for details). In the latter, each element $w_{jj^{\prime}}^{(k)}$ of the similarity matrix $\bm{W}_{k}$ measures the similarity between $z_{jk}$ and $z_{j^{\prime}k}$ for $j\neq j^{\prime}$ e.g., if $z_{jk}$ is quantitative then we set $w_{jj^{\prime}}^{(k)}=\exp(-|z_{jk}-z_{j^{\prime}k}|^{2})$, whereas if $z_{jk}$ is qualitative then $w_{jj^{\prime}}^{(k)}=1$ if $z_{jk}$ and $z_{j^{\prime}k}$ have the same categorical level (e.g., same occupation), otherwise, $w_{jj^{\prime}}^{(k)}=0$ (see also Johnson and Wichern, 1992). For completeness and reasons of parameter identifiability, the diagonals $w_{jj}^{(k)}$ are set to be zeros for all $j=1,\cdots,p$ and $k=1,\cdots,K$.
Given a set of similarity matrices $\bm{W}_{k}$ for $k=1,\cdots,K$, the Ising similarity regression model as introduced in equation (1.1) can be equivalently formulated as modeling the interaction matrix $\bm{\Theta}=(\theta_{jj^{\prime}})_{p\times p}$ via the form
$\bm{\Theta}=\sum_{j=1}^{p}\theta_{jj}\bm{\Delta}_{jj}+\sum_{k=1}^{K}\alpha_{k}\bm{W}_{k},$ (2.3)
where $\bm{\Delta}_{jj}$ is a $p\times p$ matrix with the $(j,j)$-th element being one and other elements being zeros for $j=1,\cdots,p$. This model re-parameterizes the vector $\bm{\theta}$ in equation (2.2) by a new parameter vector $\bm{\vartheta}=(\theta_{11},\cdots,\theta_{pp},\bm{\alpha}^{\top})^{\top}$ with $\bm{\alpha}=(\alpha_{1},\cdots,\alpha_{K})^{\top}$, such that the re-parameterized pmf for $\bm{y}$ can be written as
$f(\bm{u};\bm{\vartheta})=\frac{1}{Z(\bm{\vartheta})}\exp\left\{\sum_{j=1}^{p}\theta_{jj}u_{j}+\sum_{1\leq j<j^{\prime}\leq p}\left(\sum_{k=1}^{K}\alpha_{k}w_{jj^{\prime}}^{(k)}\right)u_{j}u_{j^{\prime}}\right\},$ (2.4)
where $Z(\bm{\vartheta})=\sum_{\bm{u}\in\{0,1\}^{p}}\exp\{\sum_{j=1}^{p}\theta_{jj}u_{j}+\sum_{1\leq j<j^{\prime}\leq p}(\sum_{k=1}^{K}\alpha_{k}w_{jj^{\prime}}^{(k)})u_{j}u_{j^{\prime}}\}$.
In equation (2.3), the vector of regression coefficients $\bm{\alpha}$ describes how similarities directly affect the pairwise interaction coefficients (conditional dependence relationships) of the binary responses. Note also that by utilizing similarity matrices to model the interaction matrix $\bm{\Theta}$, model (2.4) involves a substantially smaller number of parameters ($p+K$ parameters) compared to the standard Ising model in (2.2), which has $(p+1)p/2$ parameters. The proposed Ising similarity regression model is thus particularly useful when the dimension of the binary response vector $p$ is large: even when the number of similarity matrices $K$ grows at the same rate as the dimension $p$, the number of parameters in model (2.4) only grows linearly in $p$ compared to $O(p^{2})$ parameters in model (2.2).
Equation (2.3) is related to a variety of methods developed for linking a matrix parameter to a linear combination of known symmetric matrices; see, for example, Anderson (1973) and the subsequent research by Zwiernik, Uhler, and Richards (2017), and Zou et al. (2017, 2020, 2022) among others. Modeling the interaction matrix $\bm{\Theta}$ as a linear combination of the similarity matrices $\bm{W}_{k}$ for $k=1,\cdots,K$ and $\bm{\Delta}_{jj}$ for $j=1,\cdots,p$ in equation (2.3) is similar to linking a function of the covariance matrix to a linear combination of matrices (e.g., Chiu, Leonard, and Tsui, 1996; Pourahmadi, 1999; Bonat and Jørgensen, 2016), although to our knowledge this article is the first to establish such an idea for the Ising model. It is also worth noting that equation (2.3) includes the standard Ising model as a special case, by letting $\bm{W}_{1}=\bm{\Delta}_{12}+\bm{\Delta}_{21},\cdots,\bm{W}_{p-1}=\bm{\Delta}_{1p}+\bm{\Delta}_{p1},\bm{W}_{p}=\bm{\Delta}_{23}+\bm{\Delta}_{32},\cdots,\bm{W}_{p(p-1)/2}=\bm{\Delta}_{(p-1)p}+\bm{\Delta}_{p(p-1)}$, $\alpha_{1}=\theta_{12},\cdots,\alpha_{p-1}=\theta_{1p},\alpha_{p}=\theta_{23},\cdots,\alpha_{K}=\theta_{(p-1)p}$, and $K=p(p-1)/2$. As such, the Ising similarity regression model (2.4) not only allows us to estimate the interaction matrix by incorporating the similarities and/or auxiliary information, but also may be interpreted as a generalized presentation of the standard Ising model.
2.2 Estimation and Selection
Suppose we have observations of $p$-dimensional response vectors $\bm{y}_{i}=(y_{i1},\cdots,\\ y_{ip})^{\top}\in\{0,1\}^{p}$ for $i=1,\cdots,n$, where $\bm{y}_{1},\cdots,\bm{y}_{n}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}f(\cdot;\bm{\vartheta})$ follows the Ising similarity regression model in equation (2.4), and $i.i.d$ denotes independent and identically distributed. Recall the pmf $f(\cdot;\bm{\vartheta})$ involves the normalization constant $Z(\bm{\vartheta})=\sum_{\bm{u}\in\{0,1\}^{p}}\exp\{\sum_{j=1}^{p}\theta_{jj}u_{j}+\sum_{1\leq j<j^{\prime}\leq p}(\sum_{k=1}^{K}\alpha_{k}w_{jj^{\prime}}^{(k)})u_{j}u_{j^{\prime}}\}$, which is a sum of $2^{p}$ terms. As a result, maximum likelihood estimation based on $\prod_{i=1}^{n}f(\bm{y}_{i};\bm{\vartheta})$ is computationally not feasible when the dimension $p$ is large. To overcome this problem, we adapt the existing literature (e.g., Höfling and Tibshirani, 2009; Ravikumar, Wainwright, and Lafferty, 2010) and propose a pseudo-likelihood estimation approach to fit the Ising similarity regression model. Let $\bm{y}_{i\backslash j}=(y_{i1},\cdots,y_{i,j-1},y_{i,j+1},\cdots,y_{ip})^{\top}$ denote the $i$-th observed response vector without the $j$-th element, and $f_{j}(\cdot|\bm{y}_{i\backslash j};\bm{\vartheta})$ denote the corresponding conditional pmf of $y_{ij}$ given $\bm{y}_{i\backslash j}$ for $i=1,\cdots,n$ and $j=1,\cdots,p$. Then the (unregularized) pseudo-likelihood estimator can be obtained by maximizing the pseudo-likelihood function $\prod_{i=1}^{n}\prod_{j=1}^{p}f_{j}(y_{ij}|\bm{y}_{i\backslash j};\bm{\vartheta})$. In particular, the conditional pmf $f_{j}(\cdot|\bm{y}_{i\backslash j};\bm{\vartheta})$ can be derived based on $f(\cdot;\bm{\vartheta})$ in equation (2.4), and takes the simple form below,
$f_{j}(u|\bm{y}_{i\backslash j};\bm{\vartheta})$ =$=\{\mathrm{P}(y_{ij}=1|\bm{y}_{i\backslash j};\bm{\vartheta})\}^{u}\{1-\mathrm{P}(y_{ij}=1|\bm{y}_{i\backslash j};\bm{\vartheta})\}^{1-u}$
$=\frac{\exp\left\{u\left(\theta_{jj}+\sum_{k=1}^{K}\alpha_{k}\sum_{j^{\prime}\neq j}w_{jj^{\prime}}^{(k)}y_{ij^{\prime}}\right)\right\}}{1+\exp\left(\theta_{jj}+\sum_{k=1}^{K}\alpha_{k}\sum_{j^{\prime}\neq j}w_{jj^{\prime}}^{(k)}y_{ij^{\prime}}\right)},$ (2.5)
for $u\in\{0,1\}$, $i=1,\cdots,n$ and $j=1,\cdots,p$. It follows that the conditional log-odds is given by
$\log\left\{\frac{\mathrm{P}(y_{ij}=1|\bm{y}_{i\backslash j};\bm{\vartheta})}{1-\mathrm{P}(y_{ij}=1|\bm{y}_{i\backslash j};\bm{\vartheta})}\right\}=\theta_{jj}+\sum_{k=1}^{K}\alpha_{k}\sum_{j^{\prime}\neq j}w_{jj^{\prime}}^{(k)}y_{ij^{\prime}},$ (2.6)
for $i=1,\cdots,n$ and $j=1,\cdots,p$, from which we observe that the conditional pmf $f_{j}(\cdot|\bm{y}_{i\backslash j};\bm{\vartheta})$ and thus pseudo-likelihood estimation no longer involve the intractable normalization constant $Z(\bm{\vartheta})$. In fact, equation (2.6) bears a similar form to fitting a logistic regression for the conditional log-odds of $y_{ij}=1$ against the set of $K$ covariates $\{\sum_{j^{\prime}\neq j}w_{jj^{\prime}}^{(k)}y_{ij^{\prime}}=\bm{W}_{j\cdot}^{(k)\top}\bm{y}_{i}:k=1,\cdots,K\}$, with an intercept term $\theta_{jj}$ and regression coefficients $\bm{\alpha}=(\alpha_{1},\cdots,\alpha_{K})^{\top}$, where $\bm{W}_{j\cdot}^{(k)\top}\in\mathbb{R}^{1\times p}$ denotes the $j$-th row of $\bm{W}_{k}$.
We can further augment the above pseudo-likelihood estimator with a penalty to perform variable selection on the elements of $\bm{\alpha}$. This is useful in practice when there are a non-negligible number of similarity matrices available for data analysis, but only a subset of them are anticipated to be relevant in equation (2.3). For example, our real application in Section 5 considered 15 similarity matrices for the U.S. senators’ voting records data, but only a subset of these were selected to be informative for driving conditional dependence relationships in their voting patterns.
To perform simultaneous estimation and regularization on the coefficient vector $\bm{\alpha}$, we augment the log pseudo-likelihood function based on equation (2.5) with an adaptive lasso penalty (Zou, 2006), resulting in a regularized pseudo-likelihood estimator that minimizes the objective function
$-\frac{1}{np}\sum_{i=1}^{n}\sum_{j=1}^{p}\log\{f_{j}(y_{ij}|\bm{y}_{i\backslash j};\bm{\vartheta})\}+\lambda\sum_{k=1}^{K}w_{k}\left|\alpha_{k}\right|$
$=$ $-\frac{1}{np}\sum_{i=1}^{n}\sum_{j=1}^{p}\left[y_{ij}\left(\theta_{jj}+\sum_{k=1}^{K}\alpha_{k}\bm{W}_{j\cdot}^{(k)\top}\bm{y}_{i}\right)\right.$
$-\left.\log\left\{1+\exp\left(\theta_{jj}+\sum_{k=1}^{K}\alpha_{k}\bm{W}_{j\cdot}^{(k)\top}\bm{y}_{i}\right)\right\}\right]+\lambda\sum_{k=1}^{K}w_{k}\left|\alpha_{k}\right|,$ (2.7)
given a tuning parameter $\lambda>0$. Following Zou (2006) and Huang, Ma, and Zhang (2008) among others, we set the adaptive weights as $w_{k}=1/|\bar{\alpha}_{k}|$ for $k=1,\cdots,K$, where $\bar{\bm{\alpha}}=(\bar{\alpha}_{1},\cdots,\bar{\alpha}_{K})^{\top}$ denotes the unregularized pseudo-likelihood estimator i.e., the estimator of $\bm{\alpha}$ which minimizes the objective function (2.7) with $\lambda=0$. It is worth noting that $\theta_{jj}$, the $j$-specific intercepts for $j=1,\cdots,p$, are not regularized; this is similar to the literature of sparse Ising models (Guo et al., 2010; Ravikumar, Wainwright, and Lafferty, 2010) as well as other regularized regression settings in general (e.g, Hui, Müller, and Welsh, 2017b, 2018). Regarding the choice of the penalty function, in this article we focus our developments on the adaptive lasso penalty $\lambda\sum_{k=1}^{K}w_{k}\left|\alpha_{k}\right|$; see Section 6 for a discussion on alternative penalty functions. In particular, with adaptive weights there is only a single regularization parameter $\lambda$ in the penalty, and the whole objective function (2.7) remains convex for optimization. Moreover, the incorporation of adaptive weights $w_{k}$ allows for varying degrees of regularization on the coefficients $\alpha_{k}$ across $k=1,\cdots,K$, and facilitates selection consistency for a sparse coefficient vector $\bm{\alpha}$ which we theoretically examine in the next section. The details of implementing the regularized pseudo-likelihood estimation by minimizing (2.7) are discussed in Section 4.
3 Theoretical Results
In this section, we establish asymptotic properties for the regularized pseudo-likelihood estimator of the Ising similarity regression model (2.4). Since our main interest lies in estimating similarity regression coefficients $\bm{\alpha}$ that only depend on the pairwise interaction coefficients in equation (1.1), then we focus on a variant of the model with no main effects, giving rise to the criterion
$\hat{\bm{\alpha}}=\min_{\bm{\alpha}}\left\{-l(\bm{\alpha})+\lambda\sum_{k=1}^{K}w_{k}\left|\alpha_{k}\right|\right\},$ (3.8)
where $l(\bm{\alpha})=\sum_{i=1}^{n}l_{i}(\bm{\alpha})/(np)$ and $l_{i}(\bm{\alpha})=\sum_{j=1}^{p}[y_{ij}(\sum_{k=1}^{K}\alpha_{k}\bm{W}_{j\cdot}^{(k)\top}\bm{y}_{i})-\log\{1+\exp(\sum_{k=1}^{K}\alpha_{k}$ $\bm{W}_{j\cdot}^{(k)\top}\bm{y}_{i})\}]$; see Guo et al. (2010, 2015) for a similar theoretical treatment. As discussed previously, the adaptive weights $w_{k}=1/|\bar{\alpha}_{k}|$ in criterion (3.8) are constructed based on the elements in the vector of the unregularized pseudo-likelihood estimator $\bar{\bm{\alpha}}=\arg\min_{\bm{\alpha}}\{-l(\bm{\alpha})\}$. With this in mind, we will first investigate the asymptotic properties of $\bar{\bm{\alpha}}$.
Let $\bm{\alpha}^{(0)}=(\alpha^{(0)}_{1},\cdots,\alpha^{(0)}_{K})^{\top}$ denote the true value of the coefficient vector $\bm{\alpha}$, $S=\{k:\alpha^{(0)}_{k}\neq 0,\textrm{ for }k=1,\cdots,K\}$ denote the set indexing all truly non-zero coefficients, $S^{c}=\{1,\cdots,K\}\backslash S$, and let the cardinality of $S$ be denoted by $|S|=K_{0}$, where $K_{0}$ is assumed to be finite. Furthermore, we construct
$\bm{\mathcal{X}}^{(i)}=\begin{pmatrix}\bm{\mathcal{X}}^{(i,1)\top}\\ \vdots\\ \bm{\mathcal{X}}^{(i,p)\top}\end{pmatrix}=\begin{pmatrix}\bm{W}_{1\cdot}^{(1)\top}\bm{y}_{i}&\cdots&\bm{W}_{1\cdot}^{(K)\top}\bm{y}_{i}\\ \vdots&\ddots&\vdots\\ \bm{W}_{p\cdot}^{(1)\top}\bm{y}_{i}&\cdots&\bm{W}_{p\cdot}^{(K)\top}\bm{y}_{i}\end{pmatrix},\textrm{ for }i=1,\cdots,n,$ (3.9)
and define the $K\times K$ matrices $\bm{U}^{0}=E\{\sum_{i=1}^{n}\bm{\mathcal{X}}^{(i)\top}\bm{\mathcal{X}}^{(i)}/(np)\}$ and $\bm{M}^{0}=E\{-\nabla^{2}l(\bm{\alpha}^{(0)})\}$. Finally, for a generic $m\times r$ matrix $\bm{H}=(h_{tv})_{m\times r}$, and subsets of row and column indices $\mathcal{T}\subseteq\{1,\cdots,m\}$ and $\mathcal{V}\subseteq\{1,\cdots,r\}$, let $\bm{H}_{\mathcal{T},\mathcal{V}}$ denote the submatrix of $\bm{H}$ consisting of rows and columns indexed by $\mathcal{T}$ and $\mathcal{V}$, respectively, $\left\|\bm{H}\right\|_{\infty}=\max_{1\leq t\leq m}\{\sum_{v=1}^{r}|h_{tv}|\}$ denote the matrix infinity-norm of $\bm{H}$, and $\Lambda_{\min}(\bm{H})$ and $\Lambda_{\max}(\bm{H})$ denote the smallest and largest eigenvalues of $\bm{H}$, respectively, when $\bm{H}$ is a square matrix (i.e., $m=r$). We introduce the following technical conditions.
Condition 1.
There exist finite positive constants $C_{\min}$ and $C_{\max}$ such that $\Lambda_{\min}(\bm{M}^{0})\\ \geq C_{\min}$ and $\Lambda_{\max}(\bm{U}^{0})\leq C_{\max}$.
Condition 2.
There exists a finite positive constant $C_{W}$ such that $\left\|\bm{W}_{k}\right\|_{\infty}\leq C_{W}$ for all symmetric similarity matrices $\{\bm{W}_{k}:k=1,\cdots,K\}$.
By the Cauchy Interlacing Theorem (Parlett, 1980, p. 186) and Condition 1, similar conditions hold for the submatrices $\bm{M}_{S,S}^{0}$ and $\bm{U}_{S,S}^{0}$; in particular, $\Lambda_{\min}(\bm{M}_{S,S}^{0})\geq C_{\min}$ and $\Lambda_{\max}(\bm{U}_{S,S}^{0})\leq C_{\max}$. Condition 2 is similar to the conditions imposed on the similarity matrices $\bm{W}_{k}$ in the literature (see for instance Zou et al., 2017). We first establish the estimation consistency for the unregularized pseudo-likelihood estimator.
Theorem 1.
Assume Conditions 12 are satisfied. If $K\sqrt{\log(p)/n}=o(1)$ and there exists a finite positive constant $C_{\nabla}$ such that $K=o(p^{C_{\nabla}^{2}/(8C_{W}^{2})})$ as $n,p\to\infty$, then with probability tending to one it holds that
$\left\|\bar{\bm{\alpha}}-\bm{\alpha}^{(0)}\right\|_{2}\leq\bar{M}\sqrt{\frac{K\log(p)}{n}},$
for a finite positive constant $\bar{M}>4C_{\nabla}/C_{\min}$.
The proofs of Theorem 1 along with all other theoretical results are provided in the supplementary material. The above convergence rate bears a similar form to the convergence rate of many lasso-penalized estimators (e.g., Guo et al., 2010; Nghiem et al., 2022), where the factor of $K$, instead of $K_{0}$, in the numerator is expected given this is the unregularized estimator. In addition, Theorem 1 is valid under both cases of fixed $K$ and diverging $K$, so long as $K\sqrt{\log(p)/n}\to 0$ and $K/p^{C_{\nabla}^{2}/(8C_{W}^{2})}\to 0$.
To study the asymptotic properties of the regularized pseudo-likelihood estimator, we introduce two additional technical conditions.
Condition 3.
There exists a constant $C_{M}\in(0,1)$ such that $\|\bm{M}_{S^{c},S}^{0}(\bm{M}_{S,S}^{0})^{-1}\|_{\infty}\leq 1-C_{M}$.
Condition 4.
$\lambda\sqrt{n}/\{\min\{|\alpha^{(0)}_{k}|:k\in S\}\sqrt{\log(p)}\}\to 0$ and $\lambda n/\{\sqrt{K}\log(p)\}\to\infty$.
Condition 3 is commonly known as the mutual incoherence or irrepresentability condition (Hastie, Tibshirani, and Wainwright, 2015), and, together with the conditions on $\bm{M}_{S,S}^{0}$ and $\bm{U}_{S,S}^{0}$ implied by Condition 1, are similar to those assumed by Meinshausen and Bühlmann (2006) and Guo et al. (2010) among others. Condition 4 is similar to existing conditions in the literature regarding the rates of the tuning parameters for adaptive lasso regression (e.g., Wang, Li, and Leng, 2009; Hui, Müller, and Welsh, 2018).
We now state the main results of this paper for the regularized pseudo-likelihood estimator $\hat{\bm{\alpha}}=(\hat{\alpha}_{1},\cdots,\hat{\alpha}_{K})^{\top}$ of the Ising similarity regression model.
Theorem 2.
Assume Conditions 14 are satisfied. If $K\sqrt{\log(p)/n}=o(1)$ and there exists a finite positive constant $C_{\nabla}$ such that $K=o(p^{C_{\nabla}^{2}/(8C_{W}^{2})})$ as $n,p\to\infty$, then with probability tending to one it holds that
  1. (a)
    $\left\|\hat{\bm{\alpha}}-\bm{\alpha}^{(0)}\right\|_{2}\leq M\sqrt{K_{0}\log(p)/n}$ for a finite positive constant $M>4/C_{\min}$;
  2. (b)
    $\hat{\alpha}_{k}\neq 0$ for all $k\in S$ and $\hat{\alpha}_{k}=0$ for all $k\in S^{c}$.
Theorem 2(a) establishes estimation consistency of the regularized pseudo-likelihood estimator, noting its convergence rate can be faster than that of the unregularized estimator. Theorem 2(b) establishes selection consistency of the regularized pseudo-likelihood estimator i.e., it can recover the underlying sparsity pattern of $\bm{\alpha}^{(0)}$. This is attractive when the Ising similarity regression model is applied to datasets where the dimension of responses and number of similarity matrices may be large.
To conclude this section, we remark Condition 4 provides the requirements on the rate of the tuning parameter $\lambda$ based on the smallest truly non-zero coefficient $\min\{|\alpha^{(0)}_{k}|:k\in S\}$, as the basis for establishing the results in Theorem 2. For instance, if $\min\{|\alpha^{(0)}_{k}|:k\in S\}$ is bounded away from zero, then it can be verified that $\lambda=\{n/\log(p)\}^{-t}K^{v}$ for some $t\in(1/2,1)$ and $v\in[2t-3/2,2t-1]$ suffices for Condition 4 to hold. If this is relaxed and we permit $\min\{|\alpha^{(0)}_{k}|:k\in S\}$ to tend to zero at a rate satisfying $\min\{|\alpha^{(0)}_{k}|:k\in S\}\{n/\log(p)\}^{m}\to\infty$ for some $m\in(0,1/4]$, then $\lambda=O(\{\log(p)/n\}^{q})$ for some $q\in[m+1/2,3/4]$ will satisfy the requirements. In practice, since $\min\{|\alpha^{(0)}_{k}|:k\in S\}$ is unknown, then we adopt a data-driven approach to select $\lambda$ as discussed in the next section.
4 Simulation Study
We conducted a numerical study to evaluate the finite sample performance of the regularized pseudo-likelihood estimator for the Ising similarity regression model (2.4). Following Guo et al. (2010), we generated the true main effect parameters $\{\theta_{jj}^{(0)};j=1,\cdots,p\}$ independently and uniformly in the range $[-1,-0.5]\cup[0.5,1]$. For the regression coefficients of the similarity matrices, we set $K=20$ and generated $\alpha^{(0)}_{k}$ independently and uniformly from $[-0.4,-0.3]\cup[0.3,0.4]$ for $k=1,\cdots,5$ and set $\alpha^{(0)}_{k}=0$ for $k=6,\cdots,20$. This led to $K_{0}=5$ truly non-zero coefficients. Turning to the similarity matrices, we followed the framework of Zou et al. (2017) and set the diagonal elements of all the similarity matrices $\{\bm{W}_{k};k=1,\cdots,K\}$ to zero, while the off-diagonal elements were simulated as $w_{jj^{\prime}}^{(k)}=w_{j^{\prime}j}^{(k)}=\exp(-d_{jj^{\prime}}^{(k)2})$ with $d_{jj^{\prime}}^{(k)}\stackrel{{\scriptstyle i.i.d}}{{\sim}}U(p^{-1/2},p^{1/2})$ for $j,j^{\prime}=1,\cdots,p$. Based on the above settings, we then constructed the true Ising similarity regression model as in (2.4), and afterward simulated datasets consisting of $n$ multivariate binary vectors of dimension $p$ using the R package IsingSampler (Epskamp, 2020). We considered sample sizes $n\in\{50,100,200,400\}$ and different numbers of binary responses $p\in\{10,25,50,100,200\}$, and simulated 1000 datasets for each combination of $n$ and $p$.
We employed a two-step algorithm to compute the regularized pseudo-likelihood estimator. First, from the discussion below equations (2.6) – (2.7), the case of $\lambda=0$ i.e., the unregularized pseudo-likelihood estimator, can be obtained by fitting a logistic regression model with $\bm{y}_{1:n}=(y_{11},\cdots,y_{n1},\cdots,y_{1p},\cdots,y_{np})^{\top}$ as the response and $(\bm{I}_{p}\otimes\bm{1}_{n},\bm{\mathcal{X}})$ as the model matrix, where $\bm{1}_{n}$ is an $n$-dimensional vector of ones, $\otimes$ is the Kronecker product operator, and
$\bm{\mathcal{X}}=\begin{pmatrix}\bm{W}_{1\cdot}^{(1)\top}\bm{y}_{1}&\cdots&\bm{W}_{1\cdot}^{(K)\top}\bm{y}_{1}\\ \vdots&\vdots&\vdots\\ \bm{W}_{1\cdot}^{(1)\top}\bm{y}_{n}&\cdots&\bm{W}_{1\cdot}^{(K)\top}\bm{y}_{n}\\ \vdots&\vdots&\vdots\\ \bm{W}_{p\cdot}^{(1)\top}\bm{y}_{1}&\cdots&\bm{W}_{p\cdot}^{(K)\top}\bm{y}_{1}\\ \vdots&\vdots&\vdots\\ \bm{W}_{p\cdot}^{(1)\top}\bm{y}_{n}&\cdots&\bm{W}_{p\cdot}^{(K)\top}\bm{y}_{n}\\ \end{pmatrix}.$
After obtaining the unregularized estimator $\bar{\bm{\alpha}}$ in the first step, we used its elements $\bar{\alpha}_{k}$’s to construct adaptive regularization weights in (2.7) before proceeding to compute the regularized pseudo-likelihood estimator by fitting an adaptive lasso regularized logistic regression to $\bm{y}_{1:n}$ and $(\bm{I}_{p}\otimes\bm{1}_{n},\bm{\mathcal{X}})$ in the second step. We used the R package glmnet (Friedman, Hastie, and Tibshirani, 2010) to accomplish this. To select the tuning parameter $\lambda$ in (2.7), we utilized a ten-fold cross-validation approach, where the observations in each dataset were grouped at the level indexed by $i=1,\cdots,n$. That is, we split the data $\{(\bm{y}_{1},\bm{\mathcal{X}}^{(1)}),\cdots,(\bm{y}_{n},\bm{\mathcal{X}}^{(n)})\}$ randomly into ten folds, fitted the regularized logistic regression to nine of the (training) folds, and then computed the pseudo-likelihood of the fitted model on the remaining (test) fold. This was then repeated so that each fold acted as a test set once, noting that the same set of adaptive regularization weights constructed based on the unregularized estimator from the full dataset was used throughout. We then selected an optimal $\lambda$ that gave the largest mean pseudo-likelihood (averaged across the ten test sets). Note the grouping of the observations was done to preserve the dependence structure of the data, since $y_{ij}$ were independent in the index $i$ but dependent in the response index $j$ through the Ising model (see also Warton, Thibaut, and Wang, 2017). We compared the performance of our proposed regularized pseudo-likelihood estimator against two other basic approaches: the unregularized estimator based on (2.7) with $\lambda=0$ (i.e., no regularization), and the oracle estimator based on minimizing an alternative to (2.7) where $\lambda=0$ and $\alpha_{k}$ are set to be zeros for $k=6,\cdots,20$. The unregularized estimator was included to examine whether regularization could lead to better overall estimation performance for the regression coefficients and main effect parameters, while the oracle estimator was used as a benchmark since it incorporated additional information regarding the index set $S=\{1,\cdots,5\}$. For all three methods compared, we evaluated the performance of the point estimates for the regression coefficients $\alpha_{k}$ and main effect parameters $\theta_{jj}$ using the mean square error (MSE), $\textrm{MSE}_{\alpha}=(1/1000)\sum_{l=1}^{1000}\sum_{k=1}^{20}(\hat{\alpha}_{k,\{l\}}-\alpha^{(0)}_{k})^{2}/K$ and $\textrm{MSE}_{\theta}=(1/1000)\sum_{l=1}^{1000}\sum_{j=1}^{p}\\ (\hat{\theta}_{jj,\{l\}}-\theta^{(0)}_{jj})^{2}/p$, respectively, noting the MSE was scaled by the number of associated parameters, and $\hat{\alpha}_{k,\{l\}}$ and $\hat{\theta}_{jj,\{l\}}$ generically denote the estimated regression coefficients and main effect parameters, respectively, from the $l$-th simulated dataset. Additionally, we assessed model selection performance for the regularized estimators of the regression coefficients based on the true positive rate (TPR) and false positive rate (FPR), where $\textrm{TPR}=(1/1000)\sum_{l=1}^{1000}\sum_{k=1}^{5}\\ \bm{1}(\hat{\alpha}_{k,\{l\}}\neq 0)/5$, $\textrm{FPR}=(1/1000)\sum_{l=1}^{1000}\sum_{k=6}^{20}\bm{1}(\hat{\alpha}_{k,\{l\}}\neq 0)/15$, and $\bm{1}(\cdot)$ is the indicator function.
Table 1: MSE for the oracle, regularized, and unregularized estimators of the regression coefficients ($\textrm{MSE}_{\alpha}$) and main effect parameters ($\textrm{MSE}_{\theta}$) in the simulation study involving the Ising similarity regression model.
$\textrm{MSE}_{\alpha}$ $\textrm{MSE}_{\theta}$
$p$ $n$ Oracle Regularized Unregularized Oracle Regularized Unregularized
10 50 0.046 0.068 0.352 0.350 0.402 0.896
100 0.022 0.044 0.147 0.167 0.207 0.370
200 0.011 0.028 0.068 0.082 0.111 0.172
400 0.005 0.016 0.033 0.038 0.057 0.080
25 50 0.006 0.015 0.026 0.298 0.391 0.546
100 0.003 0.007 0.012 0.137 0.184 0.242
200 0.001 0.003 0.006 0.068 0.088 0.118
400 0.001 0.001 0.003 0.033 0.041 0.057
50 50 0.002 0.006 0.008 0.550 0.788 0.737
100 0.001 0.002 0.004 0.120 0.156 0.205
200 $<$0.001 0.001 0.002 0.053 0.067 0.093
400 $<$0.001 $<$0.001 0.001 0.027 0.033 0.047
100 50 0.001 0.005 0.003 2.314 1.551 2.444
100 $<$0.001 0.001 0.001 0.500 0.559 0.567
200 $<$0.001 $<$0.001 0.001 0.099 0.111 0.132
400 $<$0.001 $<$0.001 $<$0.001 0.033 0.034 0.049
200 50 $<$0.001 0.005 0.001 1.971 2.180 2.073
100 $<$0.001 $<$0.001 0.001 0.290 0.350 0.343
200 $<$0.001 $<$0.001 $<$0.001 0.073 0.076 0.102
400 $<$0.001 $<$0.001 $<$0.001 0.030 0.031 0.044
Table 1 shows that the MSEs for all three estimators of the regression coefficients $\bm{\alpha}$ decreased as the sample size $n$ and the number of responses $p$ increased. The MSEs of the main effect parameters exhibited a clear decreasing trend when $n$ increased, but no obvious patterns when $p$ increased; this is to be expected given the main effect parameter $\theta_{jj}$ was specific to each of the responses in the Ising similarity regression model for $j=1,\cdots,p$. Among the three estimators, the MSEs for the regularized estimators of the regression coefficients and main effect parameters were typically much smaller than the unregularized estimators, and in fact similar to the oracle estimators especially when $n$ and $p$ were large. This is consistent with the notion that the regularized estimators take advantage of the underlying sparsity in the model, resulting in better overall estimation performance than the unregularized estimators.
Table 2: TPR and FPR for the regularized estimators of the regression coefficients in the simulation study involving the Ising similarity regression model.
True Positive Rate (TPR) False Positive Rate (FPR)
$p\backslash n$ 50 100 200 400 50 100 200 400
10 0.337 0.490 0.654 0.824 0.167 0.202 0.225 0.235
25 0.835 0.970 0.999 1 0.218 0.218 0.197 0.167
50 0.954 0.999 1 1 0.186 0.174 0.154 0.102
100 0.976 0.998 0.999 1 0.043 0.105 0.076 0.009
200 0.925 0.998 1 1 0.043 0.058 0.005 0
Turning to the model selection performance of the regularized pseudo-likelihood estimator (Table 2), while there was slight underfitting as reflected by the comparably low TPR for the case of smallest $p=10$, both the TPR and FPR tended to one and zero, respectively, as $n$ and $p$ increased. This is consistent with Theorem 2(b), empirically demonstrating that the proposed method was able to recover the underlying sparsity in the Ising similarity regression model, while also providing empirical evidence for the use of the cross-validation approach in selecting $\lambda$.
5 Application to U.S. Senate Roll Call Voting Data
We applied the Ising similarity regression model (2.4) to roll call voting data from the U.S. Senate as part of the 117-th Congress, covering the period from 6 January 2021 to 20 May 2021 (date of data collection). Roll call voting data has previously been studied using a variety of statistical techniques including via undirected graphical models (Banerjee, El Ghaoui, and d’ Aspremont, 2008). Here, we use our model to study how voting associations between senators are associated with their similarities in various demographic attributes and social network profiles. The dataset was obtained from the U.S. Senate’s website (https://www.senate.gov/). It consists of the voting records on 199 bills by 100 senators of the 117-th Congress. The votes were coded as one for ‘Yea’ and zero for ‘Nay’, thus producing multivariate binary vectors $\bm{y}_{i}$ with elements $y_{ij}$ representing the binary vote on the $i$-th bill by the $j$-th senator. Similar to Banerjee, El Ghaoui, and d’ Aspremont (2008) and Guo et al. (2010), all bills were treated as independent. Furthermore, following Guo et al. (2010), we removed procedural votes with the ‘Yea/Nay’ proportion falling outside of the interval [0.3,0.7], as these were uncontroversial bills that did not reveal political dynamics in the U.S. Senate. The dataset also contained a small number of missing values since not all senators voted on every bill. Specifically, 1.44% of all votes were missing, and we chose to impute these with the majority vote of the senator’s party on the same bill. Note even though we used party to construct one of the similarity matrices later on in our model, we do not believe this form of imputation impacted our analysis to any great degree, given the percentage of the missing data was very small. The final dataset analyzed thus consisted of $n=138$ bills voted by $p=100$ senators.
We obtained several attributes for each senator, including their state, political party, class, and age from the U.S. Senate website, along with gender and occupation from Wikipedia. Additionally, each senator’s Twitter handle was obtained from the ‘us-senate’ GitHub project of ‘CivilServiceUSA’, and used to compute the number of tweets and number of followers for each senator. Each of these attributes was then converted into a similarity matrix $\bm{W}_{k}$, where for qualitative variables (state, party, class and gender) we set $w_{jj^{\prime}}^{(k)}=1$ if the $j$-th and $j^{\prime}$-th senators are in the same category and zero otherwise, while for quantitative variables (age, number of tweets, number of Twitter followers) we used $w_{jj^{\prime}}^{(k)}=\exp(-|z_{jk}-z_{j^{\prime}k}|^{2})$. For the occupation, given there was such a large number of categories, we considered the seven most common occupations in the sample (lawyer, executive, businessman, farmer, army, teacher, professor) and constructed a separate similarity matrix for each of these occupations. That is, $w_{jj^{\prime}}^{(k)}=1$ if the $j$-th and $j^{\prime}$-th senators have the same specific occupation and zero otherwise. Finally, we constructed a symmetric adjacency matrix summarizing the Twitter follower-followee relationship among the 100 senators, such that $w_{jj^{\prime}}^{(k)}=1$ if the $j$-th senator follows the $j^{\prime}$-th senator on Twitter or vice versa, and zero otherwise. Combined, the above resulted in a total of $K=15$ similarity matrices.
We fitted the proposed Ising similarity regression model using the regularized pseudo-likelihood estimator based on the adaptive lasso penalty to identify and quantify truly important attributes driving the voting associations between senators. As in the simulations of Section 4, the tuning parameter $\lambda$ was chosen using ten-fold cross-validation where the groupings were done at the bill level. After performing variable selection and obtained the estimated non-zero regression coefficients, we constructed 95% Wald confidence intervals for each coefficient based on standard errors obtained from the empirical sandwich covariance matrix; see Section LABEL:sec:additional_results_rollcall of the supplementary material for details of the empirical sandwich covariance matrix calculation.
Table 3 reports the estimated regression coefficients for the $K=15$ similarity matrices along with their corresponding 95% Wald confidence intervals. The estimation results for the main effect parameters can be found in the supplementary material. Using the adaptive lasso penalty, only 7 of the 15 regression coefficients were estimated to be non-zero. Of these, unsurprisingly there was statistically clear evidence that senators from the same state and/or party were more likely to vote similarly on the bills (analogous state and party effects were found in Banerjee, El Ghaoui, and d’ Aspremont, 2008; Guo et al., 2010, among others), although the former exhibited a much stronger effect. Most occupations were found to be uninformative for the conditional dependence relationships between senator voting patterns, except for businessman and lawyer, although it should be noted that the confidence interval corresponding to the effect of lawyer contained zero. The presence of a positive effect for the businessman similarity matrix could be attributed to senators who were businessmen tending to vote similarly on bills related to the economy.
Table 3: Point estimates and 95% confidence intervals (in parentheses) for the regression coefficients corresponding to the $K=15$ similarity matrices, based on fitting the Ising similarity regression model (2.4) to the U.S. Senate roll call voting data using regularized pseudo-likelihood estimation. Estimates whose corresponding confidence interval excludes zero are bolded.
Estimation of $\alpha_{k}$
State Party Class Age Gender Lawyer
2.342 0.167 0.021 0 0 0.018
(1.846,2.838) (0.153,0.181) (-0.034,0.075) 0 0 (-0.047,0.083)
Executive Businessman Farmer Army Teacher Professor
0 0.451 0 0 0 0
0 (0.018,0.884) 0 0 0 0
Tweets Followers Twitter Follower-Followee Relationship
-0.091 0 0.144
(-0.136,-0.046) 0 (0.110,0.177)
There was statistically clear evidence of a positive association between the Twitter follower-followee relationship adjacency matrix and senator voting patterns. A possible explanation for this might be that one senator who followed the other senator on Twitter had more exposure to their advocated ideologies, and hence was more likely to vote similarly. It is also possible that the senator who followed the other senator on Twitter already agreed with their ideologies in the first place, while their interactions on Twitter further reinforced such agreement, leading to positive associations in their voting patterns. The similarity in terms of senators’ popularity on Twitter, as measured by the number of followers, did not have any effect on the association between senators’ votes. Interestingly, although the effect of the similarity matrix associated with senators’ number of tweets was found to be negative, when we ran a separate analysis based on fitting an Ising similarity regression model with only this similarity matrix, its associated coefficient became positive: $\hat{\alpha}_{\textrm{Tweets}}=0.083$ with 95% confidence interval being $(0.080,0.087)$. Therefore, we conjecture the negative coefficient found in Table 3 was likely due to a large amount of information contained within this similarity matrix that could be explained by the other similarity matrices i.e., a form of matrix multicollinearity.
Figure 1: Graphs of weighted similarity/adjacency matrices $\hat{\alpha}_{k}\bm{W}_{k}$ associated with state, party and Twitter follower-followee relationship for a subset of 20 senators, where the edge width is proportional to the estimated $\hat{\alpha}_{k}$. The bottom right plot presents the weighted graph based on the estimated interaction matrix $\hat{\bm{\Theta}}$, where the edge width is proportional to its associated $\hat{\theta}_{jj^{\prime}}$ and edges between senators with different states and parties are colored red. Nodes are labeled with the state abbreviation of each senator and the color represents senator’s party.
Figure 1 presents graphs of the weighted similarity/adjacency matrices $\hat{\alpha}_{k}\bm{W}_{k}$ for selected attributes, together with the estimated interaction matrix $\hat{\bm{\Theta}}$ describing the voting associations for a subset of 20 senators, where $\hat{\bm{\Theta}}=\sum_{j=1}^{100}\hat{\theta}_{jj}\bm{\Delta}_{jj}+\sum_{k=1}^{15}\hat{\alpha}_{k}\bm{W}_{k}$ based on equation (2.3), and $\hat{\alpha}_{k}$ and $\hat{\theta}_{jj}$ are the regularized pseudo-likelihood estimators of the regression coefficients and the main effect parameters, respectively. The choice of senators to be included was made by beginning with an empty set, and sequentially adding pairs of senators who had the largest off-diagonal elements $\hat{\theta}_{jj^{\prime}}$ in $\hat{\bm{\Theta}}$ until the set contained the top 20 senators. To further improve clarity of the graph based on the estimated $\hat{\bm{\Theta}}$, we removed all edges with $\hat{\theta}_{jj^{\prime}}\leq 0.1297$ (the median of $\{\hat{\theta}_{jj^{\prime}}:j\neq j^{\prime}\}$); this explains the visually denser graph of Twitter follower-followee relationship compared to that of $\hat{\bm{\Theta}}$ in Figure 1. It can be seen that the estimated strong voting associations (i.e., large estimated values of $\theta_{jj^{\prime}}$) between senators were being driven heavily not only by the similarities in senators’ state and party affiliations, but also their relationships on Twitter as indicated by the red-colored edges in the graph of $\hat{\bm{\Theta}}$. Indeed, the fitted Ising similarity regression model yielded a log pseudo-likelihood of -1125.28, compared to -9441.01 for a null model with only the main effect parameters. This corresponds to a pseudo $R^{2}$ (Cohen et al., 2013) of approximately 0.88, indicating the selected model provided a much better fit than the null model through the inclusion of the available similarity/adjacency matrices $\bm{W}_{k}$ to explain the dependence structure.
6 Conclusion
In this article, we developed an Ising similarity regression model to study the effect of predictor information, arising in the form of similarity matrices, on conditional dependence relationships among binary responses. A computationally efficient regularized pseudo-likelihood estimator using the adaptive lasso penalty was proposed, which we demonstrated to be estimation and model selection consistent under a general setting where the number of similarity matrices $K$ and responses $p$ grow with sample size $n$. Simulations demonstrated the strong finite sample point estimation and selection performance of the proposed estimator. Applying the Ising similarity regression model to the U.S. Senate roll call voting data not only identified well-documented state and party effects in driving senator voting patterns, but also established new insights into the importance of senators’ similarities in businessman occupation and social network relationships on their voting dependence. These findings are new and differ from those obtained from the graphical structure analyses using standard Ising models (Guo et al., 2010, 2015). In particular, while graphical structure analyses aim to reveal the dependency structures among responses (senators), our method seeks to recover and quantify how different auxiliary information (senator’s attributes) influence this structure. While various methods have been proposed to address the former problem, the Ising similarity regression model is one of the first to have been developed for addressing the latter.
A logical next step would be to extend the proposed model to incorporate predictor information from both observations and responses, thus forming a sort of “double Ising similarity regression model”. In particular, this may involve accounting for spatial and/or temporal dependence between observations, where similarity matrices could then be built up based on some knowledge of spatial or temporal distance (e.g., Bonat and Jørgensen, 2016). Furthermore, establishing asymptotic normality results for the regularized pseudo-likelihood estimator would serve as a useful basis for inference, and provide some theoretical justification for our proposed empirical sandwich-based standard errors. Also, other penalty functions could be used to replace the adaptive lasso penalty for similarity selection, including the use of group or fusion penalties when the set of similarity matrices exhibits some sort of hierarchy or ordering (Hui, Müller, and Welsh, 2017a; Zhang et al., 2023). It would be useful to consider similarity matrices $\bm{W}_{k}^{(i)}$ that are heterogeneous across $i=1,\cdots,n$; this would correspond to extending the Ising similarity regression model to allow for heterogeneous interaction matrices $\bm{\Theta}^{(i)}$ for different binary response vectors $\bm{y}_{i}$ (see for instance Zou et al., 2022, in the context of covariance matrices of continuous response vectors). Finally, while this work has mainly focused on multivariate binary responses using the Ising model, similar modeling idea could be extended to quadratic exponential families (Gourieroux, Monfort, and Trognon, 1984) to allow for other response types.
Supplementary Materials
The Supplementary Material contains sample versions of Conditions 1 and 3 and proofs of the theorems, as well as additional details of application to the U.S. Senate roll call voting data.